OpenStudy (anonymous):

John's house and his school are 2km apart. One morning John walked to his school at an average speed of 30m/min for a few minutes, then ran the rest of the way at an average speed of 100m/min. If it took 34 minutes altogether, how many meters did John run?

OpenStudy (anonymous):

Let us say x = how many minutes john walked y = how many minutes john run x+y= 34 minutes so 30x + 30y = 1020 30*x + 100*y = 2km = 2000m Eliminate x: 70y = 980 y = 14 Let put y value one of the equations X + 14 = 34 x = 20

OpenStudy (anonymous):

because first variable is 20 second variable is 14

OpenStudy (anonymous):

i mean 20*30 + 14*100 = 2000

OpenStudy (anonymous):

thanks radar, bkalkan